An infinite line of charge with charge density Mt

3)A cylinder of radius a = 9.8 cm and height h = 11.1 cm is aligned with its axis along the y-axis as shown. What is the total flux ? that passes through the cylindrical surface? Enter a positive number if the net flux leaves the cylinder and a negative number if the net flux enters the cylnder.

4)Another infinite line of charge with charge density ?2 = 4.8 ?C/cm parallel to the y-axis is now added at x = 4.9 cm as shown.

What is the new value for Ex(P), the x-component of the electric field at point P?

5)What is the total flux ? that now passes through the cylindrical surface? Enter a positive number if the net flux leaves the cylinder and a negative number if the net flux enters the cylnder.

6)The initial infinite line of charge is now moved so that it is parallel to the y-axis at x = -4.9cm.

What is the new value for Ex(P), the x-component of the electric field at point P?

7)What is the total flux ? that now passes through the cylindrical surface? Enter a positive number if the net flux leaves the cylinder and a negative number if the net flux enters the cylnder.

Thanks !
Answer #1

奶 100:10-6-1600 lrı--INGA . 109 U 4100 oo 29 10

part c )

\Phi =EA=E\cdot 2\pi rh=-29361019.72\cdot 2\pi (9.8/100)(11.1/100)

\Phi =EA=2005764\, Nm^{2}/C

Note i used you value